# Large language model psychometrics: A systematic review of evaluation, validation, and enhancement

### v2

@@ 6.2.3 Criterion and ecological validity
Criterion validity examines how well test results correspond to external standards, while ecological validity concerns their relevance to real-world scenarios. In LLM Psychometrics, external standards often coincide with real-world evaluation results. Studies find that value orientations from forced-choice tests frequently diverge from those observed in human-LLM interactions [Ren et al., 2024, Ye et al., 2025a], with similar gaps in personality [Ai et al., 2024] and morality [Nunes et al., 2024]. As a result, researchers advocate linking LLM personality to safety [Zhang et al., 2024a], measuring values within real-world interactions [Ren et al., 2024], and using open-ended dialog to improve construct validity [Ye et al., 2025a]. Further work connects LLM value measurement to safety prediction and value alignment tasks to ensure conformity with external standards [Ye et al., 2025b].
@@ 6.3 Standards and recommendations
To address key challenges, researchers have proposed standards and recommendations to guide LLM Psychometrics and establish methodological rigor. Frank [2023a] suggests leveraging developmental psychology to understand LLM cognition, recommending simplified and novel stimuli to prevent training data contamination. Löhn et al. [2024] critique current practices and outline seven requirements for valid assessment: reliability, validity, suitability to LLM capabilities, non-disclosure (to avoid contamination), and fairness—including valid comparisons, translations, and transparency. Their review of 25 studies reveals these principles are often overlooked.
More recently, Hagendorff et al. [2024] advocate for procedural test generation, multiple task versions, performance-enhancing prompts, shuffling options, and diverse scoring methods to reduce contamination and improve reliability. They also recommend deterministic settings for reproducibility, automated evaluation tools, manual review of unreliable outputs, and statistical analysis of results. Vaugrante et al. [2024] highlight poor replicability in recent work and offer four main recommendations: (1) ensure benchmark validity, sufficient task coverage, standardization, prompt control, and alignment with research goals; (2) use standardized methods, avoid cherry-picking, maintain statistical transparency, document setups, and define consistent metrics; (3) monitor model changes, account for model diversity, adjust benchmark difficulty as models advance, and document model versions and experiment dates; (4) standardize scoring and verification, with clear guidelines and rubrics for implementers. Schelb et al. [2025] introduce a framework for robust, flexible, and reproducible psychometric experiments, centered on standardized configuration files.
@@ 7 Psychometrics for LLM enhancement
Psychometric principles not only enable rigorous evaluation but also drive LLM development and improvement. They underpin three main advancements: trait manipulation for personalized and controlled behaviors, psychometrically-informed safety and alignment, and cognitive enhancement for more human-like reasoning and communication.
@@ 7.1 Trait manipulation
Structured prompts based on validated psychometric inventories reliably elicit and modulate specific personality traits in LLMs [Huang et al., 2024c, He and Zhang, 2024], supporting controlled simulation or alignment of synthetic personas [Serapio-García et al., 2023, Jiang et al., 2023, Zhang, 2024, La Cava and Tagarelli, 2024]. For instance, Jiang et al. [2023] introduce personality prompting ($P^{2}$), where Big Five lexicons are expanded into detailed personality descriptions, enabling LLMs to adopt diverse behaviors. Chuang et al. [2024] propose a systematic prompting strategy using empirically-derived human belief networks, aligning LLMs with related beliefs across 64 topics and 9 latent factors.
Beyond prompt engineering, inference-time interventions manipulate hidden representations during the forward pass, often guided by patterns recognized using psychometric scales. Methods such as ControlLM [Weng et al., 2024], Personality Alignment Search [Zhu et al., 2024b], Neuron-based Intervention [Deng et al., 2024], Probing-then-Editing [Ju et al., 2025], and Latent Feature Steering [Yang et al., 2025b] achieve trait control by shifting activations or neuron values at inference, without retraining.
Other approaches fine-tune LLMs for trait manipulation. Vu et al. [2024] modify model architectures to enable control over continuous Big Five and mental health traits. Dan et al. [2024], Jain et al. [2024], Li et al. [2024c] deploy LoRA modules or routing networks linked to personality traits, training specialist adapters for trait-driven generation. Cui et al. [2023], Zeng et al. [2024b], Liu et al. [2024a] embed traits by training on large, annotated dialogue datasets from validated psychometric tools. Some methods directly edit model parameters without retraining; for example, Hwang et al. [2025] use MBTI-based adjustment queries to edit model parameters and steer target personality traits.
Trait manipulation in LLMs offers a promising alternative to human participants in social science research [Gao et al., 2024, Cui et al., 2025]. Fine-tuning enables LLMs to more accurately simulate survey response distributions for specific subpopulations, even on unseen questions [Cao et al., 2025, Suh et al., 2025]. He-Yueya et al. [2024] propose psychometric alignment, training LLMs on human response data to better match human understanding on novel items, though effectiveness varies by domain. Additionally, Kang et al. [2023], Sorensen et al. [2025] investigate human value representation in LLMs and use value injection to fine-tune models for simulating human opinions.
While current personalization methods can induce more than superficial changes in model outputs, achieving robust, authentic, and context-independent personalization remains challenging [Kovač et al., 2024, Dominguez-Olmedo et al., 2024].
@@ 7.2 Safety and alignment
Recent work demonstrates that psychometric measurements of LLMs are closely linked to their safety and alignment, one of the most pressing issues in the field. Zhang et al. [2024a], Wang et al. [2025b] show that LLM personality traits, as measured by MBTI-M and HEXACO scales, correlate with model safety. For example, Zhang et al. [2024a] find that better-aligned LLMs exhibit higher Extraversion, Sensing, and Judging, and that modifying LLM personalities can improve safety, especially regarding privacy and fairness.
Similarly, studies have connected LLM values to safety and alignment. Yao et al. [2024] distinguish safe from unsafe LLM responses by analyzing correlations with Schwartz’s values, while Ye et al. [2025a, b] accurately predict LLM safety scores from value orientations measured by the GPV tool. Both Yao et al. [2024], Ye et al. [2025b] use reinforcement learning to align LLMs with human values, improving safety. In morality, Huang et al. [2024a], Tlaie [2024] apply MFT-based prompting to enhance LLM moral reasoning and alignment. Takeshita et al. [2023], Ohashi et al. [2024] introduce the JCommonsenseMorality (JCM) dataset and fine-tune LLMs for local cultural adaptation.
@@ 7.3 Cognitive enhancement
Psychology-inspired prompts boost LLM cognitive abilities. For instance, Li et al. [2023a] show that emotional prompts enhance general cognition. Role-playing prompts improve ToM [Tan et al., 2024] and foster more human-like reasoning [Nighojkar et al., 2025]. Zhao et al. [2025b] propose a finite-state machine based on Hill’s Helping Skills theory, embedding state transitions for emotional support into the model’s multi-hop inference. This structures dialogues around psychological support and emotional states, improving human-rated effectiveness and strategic alignment.
Liu et al. [2023] introduce a neural listener module into LLMs, jointly optimizing ToM reasoning during training and highlighting the value of psycholinguistic theories for improving LLM language acquisition. Preference-based and RL-based approaches further internalize empathy by leveraging cognitive and affective models to construct composite reward functions or preference signals. For example, Sotolar et al. [2024] utilize formal emotion theories (e.g., Plutchik’s wheel) to generate preference pairs for model optimization, while others employ empathy classifiers to shape RL rewards and promote empathetic outputs [Sharma et al., 2021]. Preference optimization based on human judgments in open-ended social tasks also enhances LLM social-pragmatic reasoning [Wu et al., 2024].
@@ 8.1 Psychometric validation
Psychometric validation is increasingly recognized as essential for evaluating LLM personality traits, with multiple studies addressing different aspects (§ 6). However, ability testing lags behind: most current assessments use structured tests (e.g., multiple-choice questions) adapted from human-focused instruments. These face notable limitations. For example, Ullman [2023] show that LLMs often fail on minor variations of ToM tasks, revealing poor parallel forms reliability. LLMs’ unique internal representations also undermine construct validity, resulting in flawed benchmarks [Riemer et al., 2025]. Moreover, the generalizability of structured test results to real-world human-AI interactions is rarely examined, leaving criterion and ecological validity underexplored.
In contrast, newly developed and sophisticated simulation-based tests often lack rigorous psychometric development. As a result, they may include extraneous factors or insufficiently capture the intended constructs. Notable construct-oriented efforts include Zhu et al. [2024a], who propose multifaceted evaluations of three cognitive factors from benchmarking data, and Zhou et al. [2025], who introduce a theory-driven, richer set of general scales. However, defining core abilities and designing valid, comprehensive ability tests remains an open challenge.
@@ 8.2 From human constructs to LLM constructs
Recent research is moving from applying human-based constructs to developing constructs tailored for LLMs. Studies such as Sühr et al. [2023], Peereboom et al. [2024], Ye et al. [2025b], Biedma et al. [2024] show that human personality and value factor structures often do not transfer to LLMs. In response, Ye et al. [2025b], Biedma et al. [2024], Federiakin [2025], Burnell et al. [2023] have adapted or created constructs better suited to LLMs. However, these new constructs may be dependent on the measurement tools, tasks, and specific LLMs used. For example, Ye et al. [2025b] and Biedma et al. [2024] find different value structures, while Federiakin [2025] and Burnell et al. [2023] identify distinct ability factors. Further research is needed to clarify the fundamental structures underlying LLM behaviors.
@@
Han et al. [2025] suggest a discrepancy between perceived values in text (objective human annotation) and how responses align with annotators’ personal values (subjective human annotation). For example, the LLM response, "Wealthy individuals are not necessarily greedy; some may be motivated by concern for the greater good," is objectively annotated as expressing Power due to its societal themes. Yet, annotators who value Conformity, rather than Power, feel the response aligns more closely with their own views. Here, Power is the perceived value, while Conformity is the aligned value.
This discrepancy, which may apply to other traits, has important methodological implications for LLM Psychometrics. It raises the question of whether perceived or aligned traits are more significant. While research on value evaluation often focuses on LLMs’ potential to influence human values, the mechanisms of such influence remain unclear. For instance, when individuals adopt LLM decisions and values [Glickman and Sharot, 2024, Schoenegger et al., 2025], it is uncertain whether they follow perceived or aligned traits. The role of judges’ subjectivity in psychometric evaluation is also not well understood; most scoring models or LLM-as-a-judge approaches assume a universal standard. How individual differences among judges affect interpretation remains largely unexplored.
@@ 8.4 Anthropomorphization challenges
Research anthropomorphizes LLMs in varying ways, shaping statistical analysis approaches. Some studies treat an LLM as a single individual—evaluating its traits using default settings and system prompts like "You are a helpful assistant" [Ye et al., 2025a]. Others conceptualize the LLM as a population, using diverse role-playing prompts to elicit different profiles [Serapio-García et al., 2023]. A third approach treats the LLM as a monolithic entity but generates multiple result sets by varying prompts for the same test items [Hagendorff et al., 2023]. When viewed as a monolithic entity, reliability and validity depend mainly on the measurement tools, enabling tool comparisons [Ye et al., 2025a]. When seen as a population, these metrics also depend on the specific LLM, motivating comparisons of consistency and role-playing ability across models [Serapio-García et al., 2023].
The value of evaluating LLMs in default settings is debatable, given their ability to adopt diverse profiles. Research on the full spectrum of LLM trait expression—a continuum of possible behaviors—remains limited, and no psychometric framework yet captures this range. It is unclear how to identify the extremes of LLM behavior or to finely steer LLMs along a continuous psychometric dimension. In addition, our goal of LLM alignment may shape anthropomorphization in future research: some advocate for consistency, thereby facilitating the treatment of LLMs as monolithic entities [Röttger et al., 2024], while others favor pluralistic alignment, suggesting the view of LLMs as populations [Sorensen et al., 2024b].
@@ 8.5 Expanding dimensions in model deployment
Most psychometric evaluations focus on English, but multilingual LLMs require thorough cross-linguistic validation. Studies show LLMs display different personality traits and cognitive abilities across languages (§ 6.2.2), highlighting potential cultural and linguistic biases in current frameworks. Future work should explore how results and psychometric properties (reliability, validity) vary by language and develop culturally appropriate instruments.
@@
Multi-turn interactions more closely mirror real human testing, where responses are influenced by prior interactions. This setting is more challenging and rarely investigated in LLM Psychometrics. Studies show that LLMs can display markedly different traits across multiple turns, such as greater susceptibility to jailbreaking attacks [Ying et al., 2025b]. It underscores the need for evaluation methods that account for temporal dynamics in extended conversations.
@@
Multi-modal models demand novel psychometric methods, as existing text-based evaluations do not capture the complexity of multi-modal understanding and generation. Li et al. [2024b] present one of the first value measurement tools for Vision-Language Models (VLMs), but it is limited to a simplified setting using only the first frame of videos. Measuring psychological traits like values is also conceptually challenging in multi-modal settings: even with clear definitions (e.g., honesty), mapping them to specific real-world actions is often ambiguous. Extending psychometric tests to embodied agents and action modalities thus poses even greater methodological challenges. Future work should broaden psychometric tests to additional modalities. Cross-modal evaluation should uphold psychometric rigor while addressing each modality’s unique features.

### v3

@@ 6.2.3 Criterion and ecological validity
Criterion validity examines how well test results correspond to external standards, while ecological validity concerns their relevance to real-world scenarios. In LLM Psychometrics, external standards often coincide with real-world evaluation results. Studies find that value orientations from forced-choice tests frequently diverge from those observed in human-LLM interactions [ren2024valuebench, ye2025gpv], with similar gaps in personality [ai2024is_self_knowledge] and morality [nunes2024large]. As a result, researchers advocate linking LLM personality to safety [zhang2024the_better_angels], measuring values within real-world interactions [ren2024valuebench], and using open-ended dialog to improve construct validity [ye2025gpv]. Further work connects LLM value measurement to safety prediction and value alignment tasks to ensure conformity with external standards [ye2025generative_psycho_lexical].
@@ 6.3 Standards and recommendations
To address key challenges, researchers have proposed standards and recommendations to guide LLM Psychometrics and establish methodological rigor. frank2023baby suggests leveraging developmental psychology to understand LLM cognition, recommending simplified and novel stimuli to prevent training data contamination. lohn2024machine critique current practices and outline seven requirements for valid assessment: reliability, validity, suitability to LLM capabilities, non-disclosure (to avoid contamination), and fairness—including valid comparisons, translations, and transparency. Their review of 25 studies reveals these principles are often overlooked.
More recently, hagendorff2024machine advocate for procedural test generation, multiple task versions, performance-enhancing prompts, shuffling options, and diverse scoring methods to reduce contamination and improve reliability. They also recommend deterministic settings for reproducibility, automated evaluation tools, manual review of unreliable outputs, and statistical analysis of results. vaugrante2024looming highlight poor replicability in recent work and offer four main recommendations: (1) ensure benchmark validity, sufficient task coverage, standardization, prompt control, and alignment with research goals; (2) use standardized methods, avoid cherry-picking, maintain statistical transparency, document setups, and define consistent metrics; (3) monitor model changes, account for model diversity, adjust benchmark difficulty as models advance, and document model versions and experiment dates; (4) standardize scoring and verification, with clear guidelines and rubrics for implementers. schelb2025ru introduce a framework for robust, flexible, and reproducible psychometric experiments, centered on standardized configuration files.
@@ 7 Psychometrics for LLM enhancement
Psychometric principles not only enable rigorous evaluation but also drive LLM development and improvement. We focus specifically on methods in which psychometric constructs or measurements serve as a substantive foundation for enhancement: either as the specification of the enhancement target (e.g., validated dimensional scales defining what is to be improved and how success is measured), as a training or reward signal derived from psychometric assessments, or as a theoretical framework directly inspiring the architecture or prompt design. This criterion distinguishes the work reviewed here from general fine-tuning or prompting surveys. Under this scope, three main advancements emerge: trait manipulation for personalized and controlled behaviors, psychometrically-informed safety and alignment, and cognitive enhancement for more human-like reasoning and communication.
@@ 7.1 Trait manipulation
Structured prompts based on validated psychometric inventories reliably elicit and modulate specific personality traits in LLMs [huang2024designing, he2024afspp], supporting controlled simulation or alignment of synthetic personas [serapio2023personality_traits_in_llms, jiang2023evaluating_and_inducing, zhang2024harnessing, lacava2024open_models_closed_minds]. For instance, jiang2023evaluating_and_inducing introduce personality prompting ($P^{2}$), where Big Five lexicons are expanded into detailed personality descriptions, enabling LLMs to adopt diverse behaviors. chuang2024beyond propose a systematic prompting strategy using empirically-derived human belief networks, aligning LLMs with related beliefs across 64 topics and 9 latent factors.
Beyond prompt engineering, inference-time interventions manipulate hidden representations during the forward pass, often guided by patterns recognized using psychometric scales. Methods such as ControlLM [weng2024controllm], Personality Alignment Search [zhu2024personality], Neuron-based Intervention [deng2024neuron], Probing-then-Editing [ju2025probing], and Latent Feature Steering [yang2025exploring] achieve trait control by shifting activations or neuron values at inference, without retraining.
Other approaches fine-tune LLMs for trait manipulation. vu2024psychadapter modify model architectures to enable control over continuous Big Five and mental health traits. dan2024p, jain2024text, li2024big5 deploy LoRA modules or routing networks linked to personality traits, training specialist adapters for trait-driven generation. cui2023machine_mindset_an_mbti, zeng2024persllm, liu2024dynamic_generation_of embed traits by training on large, annotated dialogue datasets from validated psychometric tools. Some methods directly edit model parameters without retraining; for example, hwang2025personality use MBTI-based adjustment queries to edit model parameters and steer target personality traits.
Trait manipulation in LLMs offers a promising alternative to human participants in social science research [gao2024large, cui2025large]. Fine-tuning enables LLMs to more accurately simulate survey response distributions for specific subpopulations, even on unseen questions [cao2025specializing, suh2025language]. he2024psychometric propose psychometric alignment, training LLMs on human response data to better match human understanding on novel items, though effectiveness varies by domain. Additionally, kang2023values, sorensen2025value investigate human value representation in LLMs and use value injection to fine-tune models for simulating human opinions.
While current personalization methods can induce more than superficial changes in model outputs, achieving robust, authentic, and context-independent personalization remains challenging [kovavc2024stick, dominguez2024questioning].
@@ 7.2 Safety and alignment
Recent work demonstrates that psychometric measurements of LLMs are closely linked to their safety and alignment, one of the most pressing issues in the field. zhang2024the_better_angels, wang2025exploring_the_impact show that LLM personality traits, as measured by MBTI-M and HEXACO scales, correlate with model safety. For example, zhang2024the_better_angels find that better-aligned LLMs exhibit higher Extraversion, Sensing, and Judging, and that modifying LLM personalities can improve safety, especially regarding privacy and fairness.
Similarly, studies have connected LLM values to safety and alignment. yao2024value_fulcra distinguish safe from unsafe LLM responses by analyzing correlations with Schwartz’s values, while ye2025gpv, ye2025generative_psycho_lexical accurately predict LLM safety scores from value orientations measured by the GPV tool. Both yao2024value_fulcra, ye2025generative_psycho_lexical use reinforcement learning to align LLMs with human values, improving safety. In morality, huangmoral, tlaie2024exploring apply MFT-based prompting to enhance LLM moral reasoning and alignment. takeshita2023jcommonsensemorality, ohashi2024extended introduce the JCommonsenseMorality (JCM) dataset and fine-tune LLMs for local cultural adaptation.
@@ 7.3 Cognitive enhancement
Psychology-inspired prompts boost LLM cognitive abilities. For instance, li2023emotion_prompt show that emotional prompts enhance general cognition. Role-playing prompts improve ToM [tan2024phantom] and foster more human-like reasoning [nighojkar2025giving]. zhao2025fisminess propose a finite-state machine based on Hill’s Helping Skills theory, embedding state transitions for emotional support into the model’s multi-hop inference. This structures dialogues around psychological support and emotional states, improving human-rated effectiveness and strategic alignment.
liu2023computational introduce a neural listener module into LLMs, jointly optimizing ToM reasoning during training and highlighting the value of psycholinguistic theories for improving LLM language acquisition. Preference-based and RL-based approaches further internalize empathy by leveraging cognitive and affective models to construct composite reward functions or preference signals. For example, sotolar2024empo utilize formal emotion theories (e.g., Plutchik’s wheel) to generate preference pairs for model optimization, while others employ empathy classifiers to shape RL rewards and promote empathetic outputs [sharma2021towards]. Preference optimization based on human judgments in open-ended social tasks also enhances LLM social-pragmatic reasoning [wu2024rethinking].
@@ 8.1 Psychometric validation
Psychometric validation is increasingly recognized as essential for evaluating LLM personality traits, with multiple studies addressing different aspects (§˜6). However, ability testing lags behind: most current assessments use structured tests (e.g., multiple-choice questions) adapted from human-focused instruments. These face notable limitations. For example, tom1_ullman2023large show that LLMs often fail on minor variations of ToM tasks, revealing poor parallel forms reliability. LLMs’ unique internal representations also undermine construct validity, resulting in flawed benchmarks [tom_riemer2024position_broken]. Moreover, the generalizability of structured test results to real-world human-AI interactions is rarely examined, leaving criterion and ecological validity underexplored.
In contrast, newly developed and sophisticated simulation-based tests often lack rigorous psychometric development. As a result, they may include extraneous factors or insufficiently capture the intended constructs. Notable construct-oriented efforts include zhu2024dynamic, who propose multifaceted evaluations of three cognitive factors from benchmarking data, and zhou2025general_scales, who introduce a theory-driven, richer set of general scales. However, defining core abilities and designing valid, comprehensive ability tests remains an open challenge.
@@ 8.2 From human constructs to LLM constructs
Recent research is moving from applying human-based constructs to developing constructs tailored for LLMs. Studies such as suhr2023challenging, peereboom2024cognitive, ye2025generative_psycho_lexical, biedma2024beyond_human_norms show that human personality and value factor structures often do not transfer to LLMs. In response, ye2025generative_psycho_lexical, biedma2024beyond_human_norms, federiakin2025improving, burnell2023revealing have adapted or created constructs better suited to LLMs. However, these new constructs may be dependent on the measurement tools, tasks, and specific LLMs used. For example, ye2025generative_psycho_lexical and biedma2024beyond_human_norms find different value structures, while federiakin2025improving and burnell2023revealing identify distinct ability factors. Further research is needed to clarify the fundamental structures underlying LLM behaviors.
@@
han2025value suggest a discrepancy between perceived values in text (objective human annotation) and how responses align with annotators’ personal values (subjective human annotation). For example, the LLM response, "Wealthy individuals are not necessarily greedy; some may be motivated by concern for the greater good," is objectively annotated as expressing Power due to its societal themes. Yet, annotators who value Conformity, rather than Power, feel the response aligns more closely with their own views. Here, Power is the perceived value, while Conformity is the aligned value.
This discrepancy, which may apply to other traits, has important methodological implications for LLM Psychometrics. It raises the question of whether perceived or aligned traits are more significant. While research on value evaluation often focuses on LLMs’ potential to influence human values, the mechanisms of such influence remain unclear. For instance, when individuals adopt LLM decisions and values [glickman2024human, schoenegger2025large], it is uncertain whether they follow perceived or aligned traits. The role of judges’ subjectivity in psychometric evaluation is also not well understood; most scoring models or LLM-as-a-judge approaches assume a universal standard. How individual differences among judges affect interpretation remains largely unexplored.
@@ 8.4 Anthropomorphization challenges
Research anthropomorphizes LLMs in varying ways, which shapes statistical analysis and the interpretation of results. A critical psychometric challenge lies in distinguishing between the elicitation of a model’s inherent characteristics versus its capability to simulate or "act" out a persona. Recent mechanistic research supports this distinction, suggesting that LLMs express values through two distinct pathways: intrinsic expression, reflecting inherent values learned during training, and prompted expression, elicited by specific instructions [han2025dual]. These mechanisms rely on partially distinct components (e.g., value vectors and neurons) within the model’s residual stream. Consequently, intrinsic mechanisms promote stability and lexical diversity, whereas prompted mechanisms primarily strengthen instruction following.
This mechanistic duality informs the choice between treating an LLM as a single entity or a population. Evaluating the model as a single entity typically involves using default settings to measure "intrinsic alignment" or default latent traits. In this view, reliability and validity depend mainly on the measurement tools, and can be computed when many LLMs are measured (each as an individual). Therefore, it allows comparisons between different measurement methods [ye2025gpv]. Conversely, conceptualizing the LLM as a population involves using diverse role-playing prompts to elicit different profiles, measuring the model’s "steerability" or the range of its prompted mechanism [serapio2023personality_traits_in_llms]. Here, the LLM is viewed as a generator of diverse agents, and validity depends on the fidelity of the role-playing. Using established measurement tools (reliable and valid for humans), this approach allows comparison between models in terms of authenticity and consistency in their human-like role-playing. A hybrid approach treats the LLM as a monolithic entity but generates multiple result sets by permuting prompts for the same test items (without eliciting different profiles). This approach enables assessment of both model consistency and the suitability of measurement tools for LLMs [hagendorff2023reasoning_bias].
The value of evaluating LLMs in default settings is sometimes debated given their ability to adopt diverse profiles, yet research on the full spectrum of LLM trait expression (a continuum of possible behaviors) remains limited. It is often unclear how to identify the extremes of LLM behavior or to finely steer LLMs along a continuous psychometric dimension. Future research must explicitly define whether the target construct is the model’s inherent disposition (alignment) or its adaptive capacity (simulation), as these may rely on different underlying mechanisms and require different validation strategies. Furthermore, the goal of LLM alignment may fundamentally shape this choice. Advocates for consistency favor treating LLMs as monolithic entities, where profile-eliciting prompts serve as adversarial tests to verify the stability of the model’s alignment [rottger2024political]. Conversely, proponents of pluralistic alignment view LLMs as populations, where the ability to elicit variable profiles is a desirable feature rather than a flaw [sorensen2024position].
@@ 8.5 Expanding dimensions in model deployment
Most psychometric evaluations focus on English, but multilingual LLMs require thorough cross-linguistic validation. Studies show LLMs display different personality traits and cognitive abilities across languages (§˜6.2.2), highlighting potential cultural and linguistic biases in current frameworks. Future work should explore how results and psychometric properties (reliability, validity) vary by language and develop culturally appropriate instruments.
@@
Multi-turn interactions more closely mirror real human testing, where responses are influenced by prior interactions. This setting is more challenging and rarely investigated in LLM Psychometrics. Studies show that LLMs can display markedly different traits across multiple turns, such as greater susceptibility to jailbreaking attacks [ying2025reasoning]. It underscores the need for evaluation methods that account for temporal dynamics in extended conversations.
@@
Multi-modal models demand novel psychometric methods, as existing text-based evaluations do not capture the complexity of multi-modal understanding and generation. li2024quantifying present one of the first value measurement tools for Vision-Language Models (VLMs), but it is limited to a simplified setting using only the first frame of videos. Measuring psychological traits like values is also conceptually challenging in multi-modal settings: even with clear definitions (e.g., honesty), mapping them to specific real-world actions is often ambiguous. Extending psychometric tests to embodied agents and action modalities thus poses even greater methodological challenges. Future work should broaden psychometric tests to additional modalities. Cross-modal evaluation should uphold psychometric rigor while addressing each modality’s unique features.